Asset Publisher

Herschel Status Report - November 2009

Herschel Status Report - November 2009

Report for period 21 October to 17 November 2009Mission operations of the Herschel space observatory continued nominally during the reporting period, with the spacecraft and subsystems all performing as expected. The HIFI anomaly investigation has been concluded and the planned recovery activities are now expected to be formally approved on 25 November.

Spacecraft

The spacecraft is operating nominally on prime chains throughout. Preparations are well underway to conduct the first liquid helium (LHe) "Direct Liquid Content Measurement" on 25 November. This will determine the remaining amount of available liquid helium for the Herschel cooling system.

Star tracker warm pixels / autonomous switch-over
A number of specific tests were conducted over the past weeks which have resulted in a better understanding of the star tracker non-nominal behaviour: there are time-variant surplus charges on CCD pixels ("warm" pixels) and these can lead to an autonomous switch-over between the two star trackers. Investigations continue into the cause of these surplus charges, the possible evolution with time of this behaviour, and into measures that can be taken to minimise the impact these transient effects have on the Absolute Pointing Error and spatial accuracy of mapping observations.

Payload

Further instrument operational modes have been released and are being used to conduct Science Demonstration Phase and Routine Science Phase observations. Target visibility allowing, about 80% of the roughly 8800 hours of PACS & SPIRE Key Programme Time observations can now be scheduled.

As of mid-November, Herschel had about 6 days of Performance Verification Phase (PVP) activities left to do for PACS and SPIRE. Once switched back on, HIFI will require about 17 days for re-commissioning and PVP activities, some of which will only be performed in January 2010.

The cryogenic chains continue to work well and maintain the PACS and SPIRE bolometers at their operating temperatures of 300 mK.

Following intensive tests on the ground, the SPIRE on-board software has been uploaded to monitor instrument activities and autonomously recover the instrument from failures if they occur. Until there is full confidence in its correct operation the monitoring of the SPIRE spectrometer activities through the on-board control procedure, which has been in use since end August, will also continue.

HIFI recovery
The implementation of the HIFI recovery activities is being planned. While some of the activities may still be performed before Christmas, it has been decided to delay full switch-on of the instrument (including switch-on of the Local Oscillators) until early next year.

Ground Segment

Mission Operations
During the reporting period, mission operations have been conducted with the support of the ESA New Norcia ground station. Observational data stored on-board Herschel was received on ground during communication passes lasting approximately three hours.

Archiving
The ground segment is operating nominally. Following an initial period in which observation data products for science observations were delivered to users only after human-intensive individual quality checking, the availability of data products to users has now reached its Routine Phase status. This means that data products for observations with validated and released Astronomical Observation Templates (AOT) are now available to observers in the Herschel Science Archive (HSA) two days after the observation has been executed.

As more and more Key Programme observations are being scheduled and conducted, the log of executed observations and plans for upcoming observations is now made available online for the community.

Future Milestones

  • 25 November 2009: Formal approval of HIFI recovery plan
  • 25 November 2009: First "Direct Liquid Content Measurement"
  • March 2010: Herschel Observatory Routine Science Phase Readiness Review


---
Legal disclaimer
This report is based on the Herschel mission manager's report dated 17 November 2009. Please see the copyright section of the legal disclaimer (linked from the home page http://sci.esa.int) for terms of use.

Last Update: 1 September 2019
9-Mar-2021 11:26 UT

ShortUrl Portlet

Shortcut URL

https://sci.esa.int/s/w5ZgPNW

Images And Videos

Related Publications

See Also

Documentation